Sol Summary – April 20th

Crew 280 Sol Summary Report 20-04-2023

Sol: 4

Summary title: A Martian home in the Making

Author’s name: Mariona Badenas-Agustí, Commander.

Mission status: Nominal

Sol activity summary: As the Sun rose over the Martian desert, the aroma of freshly brewed coffee and warm, homemade bread filled the upper deck. Soon after, the crew gathered around the table to have a delicious breakfast and get ready for the day ahead.

During the morning, Neus Sabaté (Mission Specialist) managed to turn on her iron-based batteries in the Science Dome, while Cesca Cufí (Engineer) started building a container for her “Blob.” Meanwhile, Laia Ribas (GreenHab Officer) met with Núria Jar (Journalist) and Carla Conejo González (Executive Officer & Biologist) to discuss how to communicate her MDRS research with the public. In the observatory, Mariona Badenas-Agusti (Commander & Astronomer) aligned the telescope with Ariadna Farrés’ (Health & Safety Officer) help and was able to observe a beautiful Sunspot through the eyepiece (see Astronomy Report).

The crew reconvened for lunch, when they savored a fresh salad from the GreenHab and dehydrated scrambled eggs with vegetables. At 2PM, some crew members (Laia Ribas, Carla Conejo, Ariadna Farrés, and Mariona Badenas) started an EVA, heading to Cowboy Corner to fly a small drone, test our Midland radios, collect sand samples, and develop navigation strategies. The rest of the crew supported the EVA team and continued their research from the Hab. After returning to the base, everyone continued to work on their research.

As the Sun sets and SOL4 is drawing to a close, the crew is excited to prepare homemade pizzas with fresh tomatoes and basil from the GreenHab, complemented by a few dehydrated ingredients from the pantry. It has been a productive day, and Hypatia I looks forward to continuing their journey of exploration on the Red Planet. Building a home away from home may not be easy, but with a lot of work, effort, and (dough)-determination, anything is possible!

Look ahead plan: The crew will perform an EVA and continue their experiments throughout the day. Tomorrow, we will prepare some workshops for the weekend, including a celebration of Saint George’s special Catalan festivity.

Anomalies in work: Non-functional robotic telescope and Spirit needs a new battery wire.
Weather: Sunny, but high winds in the afternoon. Low 26F/High 55F.

Crew physical status: Nominal.

EVA: Not applicable today.

Reports to be filed: Sol Summary, EVA request, EVA Report, Astronomy Report, Operations Report, GreenHab Report, Journalist Report, Pictures of the Day.

Support Requested: (i) Replacement of stove knobs, if possible; (ii) retrieval of trash in the rear airlock.

Operations Report – April 20th

Crew 280 Operations Report 20-04-2023
SOL: 4
Name of person filing report: Cesca Cufí (Crew Engineer) and Neus Sabaté (Mission Specialist)
Non-nominal systems:
Robotic observatory. Rover Spirit.
Notes on non-nominal systems sections:
Replacement parts of Robotic observatory are scheduled to be installed during crew 261 rotation. Corroded battery wire in Rover Spirit. Replacement pending (Sergii). Rover can only be used for short distance EVA.
Spirit rover used: no
Hours: 216.7
Beginning charge: –
Ending charge: –
Currently charging: no
Opportunity rover used: yes
Hours: 121.1
Beginning charge: 100%
Ending charge: 73%
Currently charging: no
Curiosity rover used: no
Hours: 229.8
Beginning charge: –
Ending charge: –
Currently charging: no
Perseverance rover used: yes
Hours: 261.5
Beginning charge: 100%
Ending charge: 56%
Currently charging: no
General notes on rovers:
Perseverance batteries are older and discharge faster. From the station to Cowboy Corner the battery decreased down to 65%.
Summary of Hab operations:
Water use: 23 gal (from 6PM Sol3 to 6PM Sol4). Note: tank refiled yesterday, the current level of water has reached the upper section of the tank, where formula probably gives an overestimation of the water consumption.
Water (static tank): 420 gal (at 6:00PM)
Static tank pipe heater (on or off): off
Static tank heater (On or off): off
Toilet tank emptied: no
Summary of internet: used from 18:00 to 21:00 for personal communication and reports.
Summary of suits and radios:
Suits: 3,7,8 and 10.
Radios: EVA radios work nominally. Channel 1 com with the station was lost after passing by Galileo Road, driving North, as expected (from the data provided by Mission support). The EVA-Station com was maintained thanks to the crew’s own radio equipment (Midland HAM radios).
All radio and spacesuit batteries checked before EVA: ok.
Summary of GreenHab operations:
Water use: 12 gal
Heater: off
Supplemental light: off
Harvest: 21 g micro greens, 3 g mint, 116 g red tomatoes, 23 g yellow tomatoes
Summary of Science Dome operations:
Laia Ribas (Green Hab Officer): experiment.
Neus Sabaté (Mission Specialist): experiment.
Dual split: off
Summary of RAM operations: none
Summary of any observatory issues:
Robotic observatory: Not working. Replacement parts are scheduled to be installed during crew 261 rotation.
Summary of health and safety issues: none
Summary of other system’s issues:
Hab heater: it turned off about 17PM due to the high winds. Same light code as in Sol3. Crew Engineer troubleshoot the issue with the procedure given by Mission Support the previous sol (Sol3).
Hab outer layer: Hab outer layer is a little loose. When the wind is blowing, the panel is knocking the structure, constantly. The integrity of the structure is not concerned. Replacement of hab outer layer scheduled for October.
Questions, concerns and requests to Mission Support:
Spirit rover: replacement of the corroded wire
Kitchen: knobs replacement
Retrieval of trash in the rear airlock

Journalist Report – April 20th

Crew 280 Journalist Report 20-04-2023

Author: Núria Jar, Crew Journalist

The day began with no wind at the MDRS, but it was very cold outside with a temperature of only 0 Celsius degree (32ºF). The Hypatia crew members woke up warm and rested because they slept well last night. This is good news for the Crew Biologist, Carla Conejo, who is studying the impact of the simulation on the circadian rhythms of the crew. The data will show how they are already recovering and feeling rested after a stressful start of the mission. This research is essential as maintaining a healthy sleep-wake cycle for the well-being of astronauts during space missions is crucial.

Other experiments are going on their way and making good progress. The crew members are conducting the first experiments and calibrating them. One of these tests was to evaluate the feasibility of navigation strategies, which Crew Scientist, Ariadna Farrés Basiana, developed in the only extravehicular activity (EVA) of the day.

This EVA was quite long, so some crew members found themselves needing to pee during the exploration. Fortunately, this was not a problem thanks to Hypatia’s jumpsuit, which are tailored by women and can handle natural bodily functions during long EVAs like this one. The Hypatia jumpsuit has a flap in the back to make it easier to pee for women, because it is not necessary to take off the entire jumpsuit. This example highlights the importance of designing equipment and tools that take into account the specific needs and considerations of women in space.

The Hypatia I crew feels proud of being composed only by women, so it often takes advantage of any opportunity to voice its message. The crew members are proud to be an all-female crew, and they hope to inspire and encourage more women to pursue careers in STEM fields and space exploration. The four members who did the EVA took a photo with Hypatia’s flag of the mission. This is one of the moments to commemorate them later, back on Earth.

At the end of the day, Hypatia crew members are excited about the EVA request they have just submitted, pending to be approved. The purpose of this new EVA is to extract iron from rocks in order to generate batteries and demonstrate that Martian resources are sufficient to generate power and sustain a space mission. Experiments like this, led by the Crew Engineer, Neus Sabaté, are crucial towards long-term space exploration and could have implications both for future space missions and even for sustainable energy on Earth.

GreenHab Report – April 20th

Crew 280 Green Hab Report 20-04-2023

GreenHab Officer: Laia Ribas

Environmental control: Open during the day and heater during the night

Temperatures: 77.8 F

Hours of supplemental light: N/A

Daily water usage for crops: 12 gallons

Daily water usage for research and/or other purposes: N/A

Water in Blue Tank: 146 gallons

Time(s) of watering for crops: 10.00, 17:00 h

Changes to crops: N/A.

As usual, the general watering of plants and crops was done today. In the morning the door was open as it was not windy. In the afternoon the door was closed due to the wind.

Harvest: 21 g micro greens, 3 g mint, 116 g red tomatoes, 23 g yellow tomatoes

Support/supplies needed

EVA Report – April 20th

Crew 280 EVA Report 20-04-2023

EVA #4

Author: Laia Ribas (GreenHab Officer) and Carla Conejo (Executive Officer & Biologist)

Purpose of EVA: This EVA allowed us to perform different tasks:
We checked the communications between the Midland radios that we brought for a larger distance between the EVA crew and the base. We also continued to familiarize ourselves with the spacesuits and communication and safety protocols during EVAs and tried a new location to drive with the rovers.
We evaluated the feasibility of navigation strategies developed by Ariadna Farrés (Scientist & HSO).
We collected some soil and rock samples to search for organic molecules for Laia Ribas’ (GreenHab Officer) CatMart project. This procedure was recorded and documented as it was an outreach activity.
For the first time, our drone was flown by Carla Conejo (XO & Biologist; TRUST Certified) and Mariona Badenas (Commander & Astronomer).

Start time: 14:09 PM

End time: 16:15 PM

Narrative: After performing 5 minutes of depressurization in the front airlock, we drove Curiosity and Perseverance rovers to Cowboy Corner to perform four main tasks: 1) test our Midland radios; 2) develop navigation strategies away from the station; 3) collect samples for the CatMart project; 4) and maneuver our small drone.

After arriving at Galileo Road, we lost communication with the base and were only able to communicate with the Hab team with our own Midland radios. With the latter, the communication between the EVA and base teams worked well during the whole time.

When reaching our destination, we parked the rovers and walked west to collect soil and rock samples for the CatMart project and test the drone. Laia Ribas (Green Hab officer), Principal investigator of the CatMart project, together with Ariadna Farres (HSO) walked into a spot located in the north of the esplanade of Cowboy Road to obtain samples for the project and get some data for the navigation project lead by Ariadana Farres. The other crew members, Mariona Badenas (Commander & Astronomer) and Carla Conejo (Executive Officer & Biologist) performed maneuvers to perform drone tests. The tests were very successful, although there was some wind. We also took beautiful pictures with Hypatia and Earth flags.

After 1 hour, we walked back to the rovers, and we drove back to the MDRS while first stopping at Pooh’s Corner. In this spot, the communication with the base was recovered. We walked to the west to continue maneuvering the drone for 15 min to finalize our EVA driving safely to the station. Before entering the Hab, we followed the 5-minutes pressurization protocol.

Destination: Cowboy Corner

Coordinates (use UTM WSG 84): 519000 E, 4253000 N

EVA participants: 4: Laia Ribas (GreenHab Officer) (EVA Leader), Mariona Badenas (Commander & Astronomer), Ariadna Farrés (HSO), Carla Conejo (XO & Biologist)

Road(s) and routes per MDRS Map: Drive on Cow Dung Road 0110

Mode of travel: Drive on Cow Dung Road 0110 to Cowboy Corner then walk around no farther than 400 meters for sampling

Astronomy Report – April 20th



Crew 280 Astronomy Report 20-04-2023

Name: Mariona Badenas
Crew: 280 (Hypatia I)
Date: 20-04-2023

Robotic Telescope Requested (choose one) MDRS-14 MDRS-WF MLC
Objects to be Imaged this Evening: Messier 63 (NGC5055)
Images submitted with this report: None
Problems Encountered: None


Solar Features Observed: We aligned the telescope and visually observed the solar chromosphere, identifying a beautiful sunspot.
Images submitted with this report: N/A
Problems Encountered: I was not able to use the ZWO camera to image the Sun. For some reason, when I used the recommended Gain and Gamma values, I was only able to see a bright screen (see attached).

Copyright © The Mars Society. All rights reserved. | Main Site